You are the owner of a small bar, The World, and you are considering opening a bakery in a vacant area in the back of the store. You estimate that it will cost you $51,348.00 to set up the bakery and that you will generate $10,100.00 in after-tax cash flows in for the life of the store (which is expected to be 10 years.) The one concern you have is that you have limited parking; by opening the bakery you run the risk of not having enough parking for customers who enjoy your bar. You estimate that the lost sales would amount to $3,491.00 per year and that your after-tax operating margin on sales at the bar is 50.00%. If your discount rate is 14.00%, what is the NPV of opening the bakery? Answer Format: Currency: Round to: 2 decimal places. Enter Answer Here...
Answer #1

NPV = - initial cost + PV of CFs over life of contract

Cash flow each year = 10100

Lost sales = 3491

Net CF per year = 10100-3491 = $6609

Life = 10 years

Using PV formula on excel, where PMT = 6609, fv = 0, nper = 10, rate = 14%

PV = 34473.31

Initial cost = 51348

NPV = -51348 + 34473.31 = - 16874.69